Given the following equations:
G - .02G = N
N + xN = G
What is the value of x?
Format your answer as .XXXX
If you could help explain this problem to me, I would be much appreciative!

Answer by rothauserc(4718)

solve the first equation for G
G - .02G = N
.98G = N
G = N/.98
substitute for G in the second equation
N + xN = G
N + xN = N/.98
multiply both sides of = by .98
.98N +.98xN = N
divide both sides of = by N
.98 +.98x = 1
.98x = .02
x = .02/.98 = 1/49 = 0.02